Soybean meal is 18% protein; cornmeal is 9% protein. How many pounds of each should be mixed together in order to get 360-lb mixture that is 15% protein?
How many pounds of cornmeal should be in the mixture?

Answer by mananth(16946)

percent ---------------- quantity
soybean 18.00% ---------------- x pounds
Cereal 9.00% ------ 360 - x pounds
Mixture 15.00% ---------------- 360
Total 360 pounds
18.00% x + 9.00% ( 360 - x ) = 15.00% * 360
18 x + 9 ( 360 - x ) = 5400
18 x + 3240 - 9 x = 5400
18 x - 9 x = 5400 - -3240
9 x = 2160
/ 9
x = 240 pounds 18.00% soybean
120 pounds 9.00% Cereal

m.ananth@hotmail.ca